11. Write the equation of the line that passes through point (–2, –8) with a slope of 0.
A) x = –2
B) x = –8
C) y = –2
D) y = –8
Thank you very much for all you do to help those who struggle!

Found 2 solutions by Nate, checkley71:
Answer by Nate(3500)

Horizontal lines have a slope of zero.
Example: y = 2 when points: (0,2) and (1,2)
m = (2 - 2)/(1 - 0) = 0/1 = 0
All the y-values will be in the y portion for your point.
(-2, -8)
y = -8
(D)

Answer by checkley71(8403)

Y=0X+b
-8=b
D)Y=-8